can you please show me how you get the answer
find the greatest common factor
10x^8-15x^4+20x^12
thank you

Answer by mananth(16949)

10x^8-15x^4+20x^12
this can be done in two parts.
first find the GCF of the co-efficients
..
10 = 2*5
15=3*5
20=2*2*5
5 is common in all the three
so 5 is the gcf of the co-efficients.
...
x^8 = x*x*x*x*x*x*x*x
x^4=x*x*x*x
x^12= x*x*x*x*x*x*x*x*x*x*x*x
...
x^4 is common in all the three which is the GCF of the variables.
multiply GCF of co-efficients & variables
5*x^4 = 5x^4 is the GCF.
m.ananth@hotmail.ca